OpenStudy (gokuporter):

Avi bought 4 bottles of mango juice for $7.20. How much would he pay for 9 bottles of mango juice at the same price? A. $3.20 B. $5.00 C. $12.20 D. $16.20

OpenStudy (nobrainiachere):

well we would need to reverse engineer it

OpenStudy (nobrainiachere):

we need to find out how much one bottle of mango juice costs

OpenStudy (adamk):

4 bottles for $7.20 9 bottles for x Making a proportion: 4 - 7.20 9 - x Cross multiplying: 4x = 9 * 7.20 4x = 64.8 Divide by 4 x = 16.2 He will pay $16.20 for 9 bottles

OpenStudy (nobrainiachere):

\(\huge{:/}\) I was getting to that